A plane starts from rest and reaches a speed of 600m/s by the end of the runway in 10 seconds. What is the acceleration of the plan? What is the distance of the runway?



1
Expert's answer
2021-01-06T13:57:54-0500

"V_0=0;V_1=600"

"a=\frac{V_1-V_0}{t}=\frac{600}{10}=60"

"s=V_0t+\frac{at^2}{2}=0*10+\frac{60*100}{2}=3000"

Answer: a = 60 m/s2;s = 3000 m